OpenStudy (anonymous):

if f(x) = 3x + 2 and g(x) = x^2 + 1, which expression is equivalent to ( f of g)(x)? (3x + 2)(x^2 + 1) 3x^2 + 1 + 2 (3x + 2)^2 + 1 3(x^2 + 1) + 2

OpenStudy (neer2890):

f(g(x)=f(x^2+1)=3(x^2+1)+2

OpenStudy (anonymous):

thanks!